# Simulating the Visual World with Artificial Intelligence: A Roadmap

### v1

@@ 4.3 Generation 3 - Planning: Modeling the Future Evolution of Complex Systems
In the third generation of world models, planning emerges as the core capability that begins to bridge the gap between environment-centered physical simulation and agent-centered reasoning. This stage encompasses both objective physical planning, such as navigation and scene evolution based on global environmental dynamics, and the potential of subjective or agentic planning, where the model internally simulates trajectories, intentions, or counterfactual futures from a first-person perspective. In this sense, planning represents the first step toward integrating a mental world model component within a primarily physical framework, enabling the model to reason not only about how the world evolves, but also about why such evolutions may occur under different beliefs or goals.

### v2

@@ 4.3 Generation 3 - Planning: Modeling the Future Evolution of Complex Systems
In the third generation of world models [nwm, assran2025vjepa2, shin2025motionstream, zhu2025astra], planning emerges as the core capability that begins to bridge the gap between environment-centered physical simulation and agent-centered reasoning. This stage encompasses both objective physical planning, such as navigation and scene evolution based on global environmental dynamics, and the potential of subjective or agentic planning, where the model internally simulates trajectories, intentions, or counterfactual futures from a first-person perspective. In this sense, planning represents the first step toward integrating a mental world model component within a primarily physical framework, enabling the model to reason not only about how the world evolves, but also about why such evolutions may occur under different beliefs or goals.